OpenStudy (anonymous):

Maria, an experienced shipping clerk can fill a certain order in 7 hours. Tom, a new clerk needs 10 hours to do the same job. working together how long will it take them to fill the order?

OpenStudy (anonymous):

1/7 + 1/10 = (7 + 10)/ 70 = 17/70 => total time t = 70/ 17 = 4hr 7 min to fill the order.